OpenStudy (anonymous):

Solve. 14y + 20 = 4y A. –2 B. 2 C. no solution D. all real numbers

OpenStudy (dmezzullo):

A.

OpenStudy (dmezzullo):

bc

OpenStudy (anonymous):

what

OpenStudy (dmezzullo):

-14y on both sides and divide -10 into 20 and get -2

OpenStudy (anonymous):

14y + 20 = 4y 14y - 4y = -20 10y = 20 y = 2 It is B

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Find the value of x that makes the statement true. 43 = 16 + 3x A. –19 B. 29 C. 9 D. 39

OpenStudy (dmezzullo):

how would it be b?

OpenStudy (anonymous):

oops....10y = - 20....y = - 2

OpenStudy (anonymous):

my bad....it isn't

OpenStudy (dmezzullo):

Its A.

OpenStudy (anonymous):

oh Find the value of x that makes the statement true. 43 = 16 + 3x A. –19 B. 29 C. 9 D. 39

OpenStudy (anonymous):

43 = 16 + 3x 43 - 16 = 3x 27 = 3x 9 = x answer C

OpenStudy (dmezzullo):

-16 3 divided by 27 9= x

OpenStudy (anonymous):

thank u both

OpenStudy (anonymous):

Name the property for the given statement. 2(10 + 15) = 2(10) + 2(15) A. addition property of equality B. distributive property C. multiplication property of equality D. subtraction property of equality

OpenStudy (dmezzullo):

Not a problem(:

OpenStudy (anonymous):

Name the property for the given statement. 2(10 + 15) = 2(10) + 2(15) A. addition property of equality B. distributive property C. multiplication property of equality D. subtraction property of equality

OpenStudy (dmezzullo):

IDK

OpenStudy (anonymous):

dang.

OpenStudy (anonymous):

distributive :)

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

thats it thank u guys

OpenStudy (dmezzullo):

No peobs(:

OpenStudy (anonymous):

you welcome :)

OpenStudy (anonymous):

dmezzullo should get the medal because I got the first one wrong